# Hyperspectral Pansharpening: Critical Review, Tools and Future Perspectives

### v1

@@ Abstract
Hyperspectral pansharpening consists of fusing a high-resolution panchromatic band and a low-resolution hyperspectral image to obtain a new image with high resolution in both the spatial and spectral domains. These remote sensing products are valuable for a wide range of applications, driving ever growing research efforts. Nonetheless, results still do not meet application demands. In part, this comes from the technical complexity of the task: compared to multispectral pansharpening, many more bands are involved, in a spectral range only partially covered by the panchromatic component and with overwhelming noise. However, another major limiting factor is the absence of a comprehensive framework for the rapid development and accurate evaluation of new methods. This paper attempts to address this issue.
We started by designing a dataset large and diverse enough to allow reliable training (for data-driven methods) and testing of new methods. Then, we selected a set of state-of-the-art methods, following different approaches, characterized by promising performance, and reimplemented them in a single PyTorch framework. Finally, we carried out a critical comparative analysis of all methods, using the most accredited quality indicators. The analysis highlights the main limitations of current solutions in terms of spectral/spatial quality and computational efficiency, and suggests promising research directions.
To ensure full reproducibility of the results and support future research, the framework (including codes, evaluation procedures and links to the dataset) is shared on https://github.com/matciotola/hyperspectral_pansharpening_toolbox, as a single Python-based reference benchmark toolbox.
@@ I Introduction
This work tackles the above issues by providing a reliable framework that will help develop and assess new solutions for HS pansharpening. We begin by scanning the literature to identify a set of benchmark state-of-the-art (SoTA) methods. Our analysis follows an operational approach, focusing on methods that provide promising experimental results in some respect and can be faithfully replicated. After this phase, we build a new development and testing toolbox. We first assemble [^2] We use the PRISMA images, which are proprietary but can be downloaded for free with permission from the Italian Space Agency. a large dataset of high-quality PAN+HS images to enable the reliable and uniform performance assessment of all methods and to support the accurate training of DL-based methods. Then, we re-implement all the selected methods in Python (retraining the DL models) and assess their performance both at reduced resolution and full resolution using the most accredited metrics in the literature. Eventually, we provide the community with an easy-to-use toolbox comprising data, tools, and results that we hope will simplify the development of new solutions and drive progress in this field. As far as we know, this is the first attempt to provide a critical comparison for the HS pansharpening problem. A unique instance of review fully devoted to HS pansharpening is in [26] but referring to dated approaches without involving DL solutions. Moreover, in [26], the assessment has been performed using simulated data (including simulated PANs) and just at reduced resolution. Instead, in this work, real data (both HS and PAN) are exploited, and the validation relies upon both reduced resolution and full resolution procedures.
@@ II Methods
TABLE I: HS pansharpening methods.
@@
$\widetilde{\mathbf{H}}$ $[W{\times}H{\times}B]$ $R{\times}R$ upscaling of $\mathbf{H}$
@@
$\widetilde{\mathbf{X}}$ or $\widehat{\mathbf{X}}$ $[W{\times}H{\times}\ast]$ Upscaling or pansharpening of any $\mathbf{X}$
This section reviews SoTA solutions for HS pansharpening. However, rather than providing an exhaustive list of all papers published on this topic, we focus on a subset of representative methods, selected because they a) are distinctive from the methodological point of view; b) are competitive in terms of quality and/or computational efficiency; c) come with a publicly available software code (in any programming language) or are described with sufficient detail to be coded anew. All of them have been re-implemented (and assessed) in this work and eventually released as part of the proposed benchmarking toolbox.
Tab. I summarizes these methods, grouped according to the usual [2] categories of Component Substitution (CS), Multi-Resolution Analysis (MRA), Model-Based Optimization (MBO) which extends the usual Variational Optimization class, and DL. In the pansharpening literature, in addition to the fusion methods it is also custom to provide the simple expansion (EXP) of the HS (with no fusion with the PAN) obtained by an approximation of the ideal interpolator. This because such a product is a latent variable of many methods and represents a spectral benchmark. In the upcoming description, we will use the symbols recalled in Tab. II.
@@ II-A Component substitution
The methods [1, 2, 26] are based on the use of a suitable transformation that brings the data to a domain where spatial and spectral components can be easily separated. Then, the spatial component is replaced by the available PAN and the data are transformed back into the original domain. By using a linear transformation and considering the substitution of a single component, a fast pansharpening process is obtained [63]. These techniques were initially proposed for MS pansharpening and, subsequently, extended to the HS case. Preliminarily, the HS datacube $\mathbf{H}\in\mathbb{R}^{w{\times}h{\times}B}$ is spatially resized to match the target pansharpening scale using (a polynomial approximation of) the ideal interpolator [1], yielding $\widetilde{\mathbf{H}}\in\mathbb{R}^{W{\times}H{\times}B}$. The component to be replaced, $\mathbf{I}\in\mathbb{R}^{W{\times}H}$, is obtained through a simple linear combination of the upscaled bands $\{\mathbf{H}^{b}\}$ of $\widetilde{\mathbf{H}}$, through suitably defined weights $\mathbf{w}=[w_{1},\ldots,w_{B}]^{T}$:
@@
Both $\mathbf{P}$ and $\mathbf{I}$ are defined in the panchromatic domain and their difference, after histogram equalization, represents the high-frequency details missing in the original image. Therefore, the pansharpened bands $\{\widehat{\mathbf{H}}^{b}\}$ of $\widehat{\mathbf{H}}\in\mathbb{R}^{W{\times}H{\times}B}$ are obtained by injecting these details in the upscaled HS bands, weighted by suitable coefficients, the injection gains $\mathbf{g}=[g_{1},\ldots,g_{B}]$
@@ II-B Multiresolution analysis
Methods based on Multi-Resolution Analysis (MRA) [1, 2, 26] are formally very similar to CS methods, as they also involve the injection of spatial details in HS upscaled bands:
@@ II-D Deep learning
As already said, DL is by far the most popular approach for MS and HS pansharpening, these days. Following our criteria, we selected seven SoTA methods for our toolbox, five of them based on supervised learning [34, 38, 35, 36, 62], and two [37, 50] on unsupervised learning. Due to the lack of GT, supervised DL models are trained following the classic Wald’s protocol [79], as suggested in [31]. This consists of low-pass filtering and decimating (in both directions) both PAN and HS by the same factor $R$, equal to the PAN-MS resolution ratio. These downscaled images are then used as synthetic PAN and HS input for training the pansharpening network, using the original HS data as GT. The network trained on reduced resolution data is then used to pansharpen full resolution data, relying on a scale invariance assumption. This latter hypothesis, however, is rather shaky, which is why unsupervised DL methods are appearing with increasing frequency lately.
@@ II-D1 HyperPNN
It is worth underlining that the HyperPNN network, like other networks described later on, is designed to work with a fixed number of bands. This is a major limitation in HS pansharpening where the number of available bands changes from sensor to sensor and, due to noise, from image to image. In fact, to deal with the three images of the dataset (Tab. IV) the authors had to train three different image-dependent networks.
@@ II-D3 DHP-DARN
This method [35] relies on two key elements, the use of a deep hyperspectral prior (DHP) model aimed at improving the preliminary upscaling of $\mathbf{H}$, and the use of the dual-attention residual network (DARN). Deep image priors (DIP) [81] are called upon to make up for the scarcity of data typical of many problems. The idea is that, lacking sufficient training data, the input to the network should be as close as possible to the expected output. In our case, the input, $\widetilde{\mathbf{H}}$, should be close to the expected result, $\widehat{\mathbf{H}}$, or, at least, spectrally consistent with the original $\mathbf{H}$. The overall effect is a sort of prior regularization that prevents possible generalization issues. Unlike other upsampling options, that work band-wise regardless of spectral dependencies, the DHP module is tuned online on the very same target image to guarantee that the upscaled $\widetilde{\mathbf{H}}$, when degraded, will return to $\mathbf{H}$. Then, the actual fusion process is carried out through the main network, denoted by the function $f_{\rm DARN}$, which consists of a sequence of three subnets by-passed by a global skip connection,
@@ II-D4 DIP-HyperKite
Another approach based on the deep image prior, called DIP-HyperKite, has been proposed in [36]. In this case, the generated prior image $\widetilde{\mathbf{H}}$ is forced to be consistent not only with $\mathbf{H}$ but also with $\mathbf{P}$. This is achieved through an additional loss term that compares $\mathbf{P}$ to a weighted average of $\widetilde{\mathbf{H}}$ along the spectral dimension, where the weights are also learned. A residual learning scheme is used, like in DARN. However, an innovative architecture is proposed here, a sort of inverse U-Net [82] where pooling and unpooling operations are exchanged, with the latter working on the “encoding” side and the former moved to the “decoding” part. By doing so, in the central part of the network, the spatial resolution increases up to eight times in both directions compared to the target resolution. This overcomplete representation is used because the residue to be predicted, $\widehat{\mathbf{H}}-\widetilde{\mathbf{H}}$, is mostly concentrated in the higher frequencies. A spatial expansion allows to widening of the frequency domain beyond the limits set by the PAN resolution, increasing the network’s ability to synthesize high-frequency spatial details. It goes without saying that the computational complexity increases considerably, both in the training and inference phases.
@@ II-D5 HyperDSNet
A partial response to these shortcomings is given by DHP-DARN and DIP-HyperKite, which use deep image priors to balance weakly trained networks. The more natural solution, however, is to use unsupervised training schemes [46, 83, 84, 47], which exploit only original full-resolution data to train the network with no need for GT. In this case, specific loss functions must be defined and carefully designed to drive the network towards the desired behavior. These losses comprise at least two terms, referred to as spectral (${\cal L}_{\lambda}$) and spatial (${\cal L}_{S}$) consistency loss terms,
@@ II-D7 PCA-Z-PNN
$\mathbf{W}$ $=$ ${\rm eig}(\widetilde{\mathbf{H}}^{T}\widetilde{\mathbf{H}}),$
@@
where, $\widetilde{\mathbf{H}}$ is zero-meaned and reshaped as a $WH{\times}B$ matrix, $\mathbf{W}$ is the $B{\times}B$ matrix whose columns are the eigenvectors of $\widetilde{\mathbf{H}}^{T}\widetilde{\mathbf{H}}$, $f_{\rm Z{-}PNN}$ is the pansharpening function, and $(\cdot)^{T}$ and $(\cdot)^{-1}$ denote transpose and inverse, respectively.
@@ III Quality assessment
In the rest of this Section, we briefly review the quality indices considered in our toolbox, also summarized in Tab. III. However, it is worth emphasizing once again that quality assessment in pansharpening is an ill-posed problem, with no simple solution, and is still the subject of intense research. Both the reduced resolution and full resolution approaches have advantages and disadvantages. A good practice, consistently followed in the literature, is to use a wide range of indices and to always integrate the numerical results with a critical visual inspection of the images by experts.
@@ III-B1 Khan’s spectral distortion index
where $\widehat{\mathbf{H}}_{\downarrow}$ indicates the MTF-based low-pass filtered and decimated version of the pansharpened image $\widehat{\mathbf{H}}$ while $\mathbf{H}$ is the original HS. Since the index is defined as $1-Q2^{n}$, the optimum value is 0 (hence the “distortion” name), and it is obtained when the downscaled version of $\widehat{\mathbf{H}}$ matches the original HS in terms of first-and second-order statistics in the multiband space.

### v2

@@ Abstract
Hyperspectral pansharpening consists of fusing a high-resolution panchromatic band and a low-resolution hyperspectral image to obtain a new image with high resolution in both the spatial and spectral domains. These remote sensing products are valuable for a wide range of applications, driving ever growing research efforts. Nonetheless, results still do not meet application demands. In part, this comes from the technical complexity of the task: compared to multispectral pansharpening, many more bands are involved, in a spectral range only partially covered by the panchromatic component and with overwhelming noise. However, another major limiting factor is the absence of a comprehensive framework for the rapid development and accurate evaluation of new methods. This paper attempts to address this issue. We started by designing a dataset large and diverse enough to allow reliable training (for data-driven methods) and testing of new methods. Then, we selected a set of state-of-the-art methods, following different approaches, characterized by promising performance, and reimplemented them in a single PyTorch framework. Finally, we carried out a critical comparative analysis of all methods, using the most accredited quality indicators. The analysis highlights the main limitations of current solutions in terms of spectral/spatial quality and computational efficiency, and suggests promising research directions. To ensure full reproducibility of the results and support future research, the framework (comprising methods and assessment codes, and dataset setup procedures) is shared on https://github.com/matciotola/hyperspectral_pansharpening_toolbox, as a single Python-based reference benchmark toolbox.
@@ I Introduction
This work tackles the above issues by providing a reliable framework that will help develop and assess new solutions for HS pansharpening. We begin by scanning the literature to identify a set of benchmark state-of-the-art (SoTA) methods. Our analysis follows an operational approach, focusing on methods that provide promising experimental results in some respect and can be faithfully replicated. After this phase, we build a new development and testing toolbox. We first assemble [^2] We use the PRISMA images, which are proprietary but can be downloaded for free with permission from the Italian Space Agency. a large dataset of high-quality PAN+HS images to enable the reliable and uniform performance assessment of all methods and to support the accurate training of DL-based methods. Then, we re-implement all the selected methods in a unique Python-based framework. This is achieved by either using the original codes as baseline, when available, or following the authors descriptions and guidelines. Among these, the DL-based ones are all retrained on the same suitably designed wide and rich dataset. Finally, we extensively compare the selected methods on the test datasets at both original resolution (or full resolution, FR), with no-reference quality indexes due to the lack of ground truth, and in a reduced resolution (RR) space using reference-based accuracy indexes that quantify the error directly. The accuracy indexes, selected among the most credited ones in the literature, are also provided as part of the toolbox. Hereinafter, the terms FR and RR data are always referred to the input given to the pansharpening process and can also be also termed real and simulated data, respectively, as the latter indicates data obtained through a suitable degradation processing of the former. Eventually, we provide the community with an easy-to-use toolbox comprising data, tools, and results that we hope will simplify the development of new solutions and drive progress in this field. As far as we know, this is the first attempt to provide a critical comparison for the HS pansharpening problem. A unique instance of review fully devoted to HS pansharpening is in [26] but referring to dated approaches without involving DL solutions. Moreover, in [26], the assessment has been performed using simulated data (including simulated PANs) and just at reduced resolution. Instead, in this work, real data (both HS and PAN) are exploited, and the validation relies upon both reduced resolution and full resolution procedures.
@@ II Methods
TABLE I: Benchmark HS pansharpening methods.
@@
$\widetilde{\mathbf{H}}$ $[W{\times}H{\times}B]$ $R{\times}R$ upsampling of $\mathbf{H}$
@@
$\widetilde{\mathbf{X}}$ or $\widehat{\mathbf{X}}$ $[W{\times}H{\times}\ast]$ Upsampling or pansharpening of any $\mathbf{X}$
This section reviews SoTA solutions for HS pansharpening. Though described or mentioned, not all reviewed methods are actually part of the experimental framework. Our critical review, in fact, aims to achieve a fair comparison where the involved methods are developed in the same framework and trained/optimized on the same datasets to prevent experimental biases. On the other hand, an overall limit to the number of solutions to develop and assess is also necessary to achieve a handy and trustworthy benchmark. Eventually, the experimental part will focus on a subset of representative methods, selected because they: a) are distinctive from the methodological point of view; b) are competitive in terms of quality and/or computational efficiency; c) come with a publicly available software code (in any programming language) or are described with sufficient detail to be faithfully coded anew.
The methods actually implemented and assessed, released as part of the proposed benchmarking toolbox, are gathered in Tab. I, where they are grouped according to the categories proposed in [2], i.e: Component Substitution (CS), Multi-Resolution Analysis (MRA), Model-Based Optimization (MBO) which extends the usual Variational Optimization class, and DL. In the pansharpening literature, in addition to the fusion methods it is also custom to provide the simple expansion (EXP) of the HS (with no fusion with the PAN) obtained by an approximation of the ideal interpolator. This because such a product is a latent variable of many methods and represents a spectral benchmark. For the sake of readability, the following sectioning matches the categories given in Tab. I and paragraphs correspond to the associated methods, though variants or alternative solutions out of benchmark are also mentioned or briefly detailed. In the upcoming description, we will use the symbols recalled in Tab. II.
@@ II-A Component substitution
CS methods are based on the use of a suitable transformation that brings the data to a domain where spatial and spectral components can be easily separated. Then, the spatial component is replaced by the available PAN and the data are transformed back into the original domain. By using a linear transformation and considering the substitution of a single component, a fast pansharpening process is obtained [63]. These techniques were initially proposed for MS pansharpening and, subsequently, extended to the HS case. Preliminarily, the HS datacube $\mathbf{H}\in\mathbb{R}^{w{\times}h{\times}B}$ is spatially resized to match the target pansharpening scale using (a polynomial approximation of) the ideal interpolator [1], yielding $\widetilde{\mathbf{H}}\in\mathbb{R}^{W{\times}H{\times}B}$. The component to be replaced, $\mathbf{I}\in\mathbb{R}^{W{\times}H}$, is obtained through a simple linear combination of the upsampled bands $\{\mathbf{H}^{b}\}$ of $\widetilde{\mathbf{H}}$, through suitably defined weights $\mathbf{w}=[w_{1},\ldots,w_{B}]^{T}$:
@@
Both $\mathbf{P}$ and $\mathbf{I}$ are defined in the panchromatic domain and their difference, after histogram equalization, represents the high-frequency details missing in the original image. Therefore, the pansharpened bands $\{\widehat{\mathbf{H}}^{b}\}$ of $\widehat{\mathbf{H}}\in\mathbb{R}^{W{\times}H{\times}B}$ are obtained by injecting these details in the upsampled HS bands, weighted by suitable coefficients, the injection gains $\mathbf{g}=[g_{1},\ldots,g_{B}]$
@@ II-B Multiresolution analysis
Methods based on Multi-Resolution Analysis (MRA) [1, 2, 26] are formally very similar to CS methods, as they also involve the injection of spatial details in HS upsampled bands:
@@ II-D Deep learning
As already said, DL is by far the most popular approach for MS and HS pansharpening, these days. Following our criteria, we have selected seven SoTA methods for our toolbox, five of them based on supervised learning [34, 38, 35, 36, 62], and two [37, 50] on unsupervised learning. Due to the lack of GT, supervised DL models are trained following the classic Wald’s protocol [79], as suggested in [31]. This consists of low-pass filtering and decimating (in both directions) both PAN and HS by the same factor $R$, equal to the PAN-MS resolution ratio. These downsampled images are then used as synthetic PAN and HS input for training the pansharpening network, using the original HS data as GT. The network trained on reduced resolution data is then used to pansharpen full resolution data, relying on a scale invariance assumption. This latter hypothesis, however, is rather shaky, which is why unsupervised DL methods are appearing with increasing frequency lately.
@@ II-D1 HyperPNN
It is worth underlining that the HyperPNN network, like other networks described later on, is designed to work with a fixed number of bands. This is a major limitation in HS pansharpening where the number of available bands changes from sensor to sensor and, due to noise, from image to image. In fact, to deal with the three images of the dataset (Tab. V) the authors had to train three different image-dependent networks.
@@ II-D3 DHP-DARN
This method [35] relies on two key elements, the use of a deep hyperspectral prior (DHP) model aimed at improving the preliminary upsampling of $\mathbf{H}$, and the use of the dual-attention residual network (DARN). DIP [81] are called upon to make up for the scarcity of data typical of many problems. The idea is that, lacking sufficient training data, the input to the network should be as close as possible to the expected output. In our case, the input, $\widetilde{\mathbf{H}}$, should be close to the expected result, $\widehat{\mathbf{H}}$, or, at least, spectrally consistent with the original $\mathbf{H}$. The overall effect is a sort of prior regularization that prevents possible generalization issues. Unlike other upsampling options, that work band-wise regardless of spectral dependencies, the DHP module is tuned online on the very same target image to guarantee that the upsampling $\widetilde{\mathbf{H}}$, when degraded, will return to $\mathbf{H}$. Then, the actual fusion process is carried out through the main network, denoted by the function $f_{\rm DARN}$, which consists of a sequence of three subnets by-passed by a global skip connection,
@@ II-D4 DIP-HyperKite
Another approach based on the DIP, called DIP-HyperKite, has been proposed in [36]. In this case, the generated prior image $\widetilde{\mathbf{H}}$ is forced to be consistent not only with $\mathbf{H}$ but also with $\mathbf{P}$. This is achieved through an additional loss term that compares $\mathbf{P}$ to a weighted average of $\widetilde{\mathbf{H}}$ along the spectral dimension, where the weights are also learned. A residual learning scheme is used, like in DARN. However, an innovative architecture is proposed here, a sort of inverse U-Net [82] where pooling and unpooling operations are exchanged, with the latter working on the “encoding” side and the former moved to the “decoding” part. By doing so, in the central part of the network, the spatial resolution increases up to eight times in both directions compared to the target resolution. This overcomplete representation is used because the residue to be predicted, $\widehat{\mathbf{H}}-\widetilde{\mathbf{H}}$, is mostly concentrated in the higher frequencies. A spatial expansion enables the frequency domain to extend beyond the limitations imposed by PAN resolution, increasing the network’s ability to synthesize high-frequency spatial details. It goes without saying that the computational complexity increases considerably, both in the training and inference phases.
@@ II-D5 HyperDSNet
A partial response to these shortcomings is given by DHP-DARN and DIP-HyperKite, which use DIPs to balance weakly trained networks. The more natural solution, however, is to use unsupervised training schemes [46, 83, 84, 47], which exploit only original full-resolution data to train the network with no need for GT. In this case, specific loss functions must be defined and carefully designed to drive the network towards the desired behavior. These losses comprise at least two terms, referred to as spectral (${\cal L}_{\lambda}$) and spatial (${\cal L}_{S}$) consistency loss terms,
@@ II-D7 PCA-Z-PNN
${[\mathbf{W},\sim,\sim]}$ $=$ ${\rm SVD}(\widetilde{\mathbf{H}}\widetilde{\mathbf{H}}^{T}),$
@@
where $\widetilde{\mathbf{H}}$ is zero-meaned and reshaped as a $WH{\times}B$ matrix, $\mathbf{W}$ is the $B{\times}B$ matrix whose columns are the eigenvectors of $\widetilde{\mathbf{H}}\widetilde{\mathbf{H}}^{T}$ obtained via SVD decomposition, $f_{\rm Z{-}PNN}$ is the pansharpening function, and $(\cdot)^{T}$ and $(\cdot)^{-1}$ denote transpose and inverse, respectively.
@@
II-E Beyond the benchmark
In addition to the above-reviewed methods, there are recent DL solutions not enclosed in the benchmark that are also worth mentioning. HyperTransformer [87] is one such method that, leveraging recent advances in the computer vision domain, involves the use of transformers [88, 89]. Specifically, it makes use of a self-attention mechanism that allows the training of the attention modules by means of a suitable loss that combines a $\ell_{1}$-norm with a perceptual term [90]. Other transformer-based approaches include MDTP-Net [91], a multi-scale pansharpening network that uses multi-head attention mechanisms and dual Gaussian-Laplacian pyramids, and MFT-GAN [92], which leverages transformers in an unsupervised adversarial training framework.
Instead, DMOEAD [93] explores multi-task learning in the context of HS pansharpening. In this work a framework is proposed which enhances the performance and generalization of HS pansharpening by incorporating the classification task into the learning process.
Another interesting solution is RE-RANet (Ratio Estimation and Residual Attention Network) [49], where the use of a residual attention mechanism is combined with the estimation (via neural networks) of the high-low resolution ratio image. Such a ratio is estimated over the PAN component and then used as a starting point for the estimation of that of each HS band. The estimation is performed employing a suitable network trained with an unsupervised loss.
An approach similar to PCA-Z-PNN is instead proposed in [94] and called PLRDiff. Likewise PCA-Z-PNN, it leverages the low-rank structure in the spectral domain that allows a dimensionality reduction. Besides, it makes use of a pre-trained diffusion model originally designed for the change detection task [95] and suitably adapted to the pansharpening of the reduced set of transformed spectral bands.
@@ III Quality assessment
In the rest of this Section, we briefly review the quality indices considered in our toolbox, also summarized in Tab. III.
However, it is worth emphasizing that any quality assessment in pansharpening is an open and still debated problem. On the one hand, many researchers [83, 99, 100, 48] have questioned the validity of scale invariance, which is a key assumption in low-resolution evaluation. On the other hand, no-reference metrics are often adapted from multispectral imaging, without properly addressing the specific challenges of HS images. Additionally, there is still no widespread agreement on the reliability of these full-resolution metrics [97, 101, 102, 103, 86, 98, 104].
Both the reduced resolution and full resolution approaches have advantages and disadvantages. A good practice, consistently followed in the literature, is to use a wide range of indices and to always integrate the numerical results with a critical visual inspection of the images by experts.
@@ III-B1 Khan’s spectral distortion index
where $\widehat{\mathbf{H}}_{\downarrow}$ indicates the MTF-based low-pass filtered and decimated version of the pansharpened image $\widehat{\mathbf{H}}$ while $\mathbf{H}$ is the original HS. Since the index is defined as $1-Q2^{n}$, the optimum value is 0 (hence the “distortion” name), and it is obtained when the downsampled version of $\widehat{\mathbf{H}}$ matches the original HS in terms of first-and second-order statistics in the multiband space.